reds:

Please help... Find the volume of the space inside the right square prism which is not occupied by the two congruent right square pyramids. The height (h) of each square pyramid is 12 cm and the length of the side of the square base is 8 cm.

reds:

1 attachment
HeartTricker:

Hi, do you know anything about volume?

reds:

@hearttricker yeah I'm just a bit lost in the problem do you know how to help?

HeartTricker:

Yes the Volume is \[\frac{ 1 }{3 } a^2h\] so you do \[\frac{ 1 }{ 3 }x 8 ^2x 12= ?\]

HeartTricker:

? would be your answer :)

AZ:

I disagree with HeartTricker

AZ:

We have a right square prism And there are two right square pyramids inside We want to find the volume NOT occupied by the pyramids That means if we find the total volume of the prism and we need to subtract the volume of the two pyramids That would get us the final answer which is the volume of the remaining space in the pyramid

AZ:

Volume of the prism is going to be Volume = Base * height Base is the area of the base. Since it's a square base, the area of the base is side^2 and each side is 8 cm so the area of the base is 8 * 8 and the height of the prism is 12 + 12 = 24 So the volume of the prism is 8 * 8 * 24 = ??

AZ:

The volume of the square pyramid is \(\dfrac{1}{3} s^2 h\) the side is 8 and the height of each pyramid is 12 so find the volume of one pyramid and then multiply by 2 because we need the volume of two pyramids

AZ:

To get to your final answer, we need to do (volume of prism) - (volume of two pyramids) = ??

AZ:

If that's too much for you to understand all at once, I can go over it with you step by step

reds:

I am working on it but I think that would be nice if you would like to go over step by step

reds:

1 attachment
reds:

am I somewhat on the right track?

ItzAgar:

I don't know the answer, but I hope this could help you at least. https://www.chino.k12.ca.us/cms/lib/CA01902308/Centricity/Domain/4926/12-5_Volumes_of_Pyramids_and_Cones.pdf

AZ:

@reds wrote:
am I somewhat on the right track?
The height of the prism is 24 not 12! If you look in the image, you see that each 'h' is 12 feet and the prism is made up of 2 * h So the volume of the prism is 8 * 8 * 24 = ?? And for the area of the pyramids you forgot to multiply by 1/3 The height of the pyramid is 12 Volume of pyramid = 1/3 * B * h = 1/3 * 64 * 12 and then you multiply by 2 to get the area of 2 pyramids

AZ:

and then you subtract final answer = (prism volume) minus (2 pyramids volume)